1. Sides of triangles are given below. Determine which of them are right triangles? In case of a right triangle, write the length of its hypotenuse.

(i) 7 cm, 24 cm, 25 cm
(ii) 3 cm, 8 cm, 6 cm
(iii) 50 cm, 80 cm, 100 cm
(iv) 13 cm, 12 cm, 5 cm

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
(i) Given, sides of the triangle are 7 cm, 24 cm, and 25 cm. Squaring the lengths of the sides of the, we will get 49, 576, and 625. 49 + 576 = 625 (7)2 + (24)2 = (25)2 Therefore, the above equation satisfies, Pythagoras theorem. Hence, it is right angled triangle. Length of Hypotenuse = 25 cm (ii) Given, sides of the triangle are 3 cm, 8 cm, and 6 cm. Squaring the lengths of these sides, we will get 9, 64, and 36. Clearly, 9 + 36 ≠ 64 Or, 32 + 62 ≠ 82 Therefore, the sum of the squares of the lengths of two sides is not equal to the square of the length of the hypotenuse. Hence, the given triangle does not satisfies Pythagoras theorem. (iii) Given, sides of triangle’s are 50 cm, 80 cm, and 100 cm. Squaring the lengths of these sides, we will get 2500, 6400, and 10000. However, 2500 + 6400 ≠ 10000 Or, 502 + 802 ≠ 1002 As you can see, the sum of the squares of the lengths of two sides is not equal to the square of the length of the third side. Therefore, the given triangle does not satisfies Pythagoras theorem. Hence, it is not a right triangle. (iv) Given, sides are 13 cm, 12 cm, and 5 cm. Squaring the lengths of these sides, we will get 169, 144, and 25. Thus, 144 +25 = 169 Or, 122 + 52 = 132 The sides of the given triangle are satisfying Pythagoras theorem. Therefore, it is a right triangle. Hence, length of the hypotenuse of this triangle is 13 cm.